# Weight-Space Linear Recurrent Neural Networks

### v1

@@ Abstract
We introduce WARP (Weight-space Adaptive Recurrent Prediction), a simple yet powerful framework that unifies weight-space learning with linear recurrence to redefine sequence modeling. Unlike conventional recurrent neural networks (RNNs) which collapse temporal dynamics into fixed-dimensional hidden states, WARP explicitly parametrizes the hidden state as the weights of a distinct root neural network. This formulation promotes higher-resolution memory, gradient-free adaptation at test-time, and seamless integration of domain-specific physical priors. Empirical validation shows that WARP matches or surpasses state-of-the-art baselines on diverse classification tasks, spanning synthetic benchmarks to real-world datasets. Furthermore, extensive experiments across sequential image completion, dynamical system reconstruction, and multivariate time series forecasting demonstrate its expressiveness and generalization capabilities. Critically, WARP’s weight trajectories offer valuable insights into the model’s inner workings. Ablation studies confirm the architectural necessity of key components, solidifying weight-space linear RNNs as a transformative paradigm for adaptive machine intelligence.
@@ 1 Introduction
Deep sequence models consistently demonstrate limitations in generating trajectories that diverge from their training distribution Alijani et al. (2024); Hataya et al. (2024); Graham et al. (2023). For instance, subsets of neural ODE Chen et al. (2018) parameters necessitate adaptation via gradient descent to maintain performance on out-of-distribution (OoD) sequences Kassaï Koupaï et al. (2024); Nzoyem et al. (2025a). While powerful, their inherent gradient calculation cost has recently catalyzed research into gradient-free alternatives Serrano et al. (2024); Morel et al. (2025); Hemmer and Durstewitz (2025), essentially trading accuracy for computational efficiency. In an effort to preserve both while achieving true zero-shot adaptation, we investigate two powerful emerging paradigms: weight-space learning and linear recurrence Rumelhart et al. (1986).
@@
Weight-space learning — the paradigm that treats the weights and biases of a neural network as data points for another learning system Schürholt et al. (2024) — offers unprecedented potential for extracting properties of a trained model solely from its “weights” [^1] Following the convention from Zhou et al. (2023), we refer to the learnable parameters of the processed network as ‘weights’ (or ‘weight-space’ to indicate the space they belong to) and those of the higher-level learning system (the neural functional) as simply ‘parameters’.. Applications span from predicting generalization error Unterthiner et al. (2020) and recovering training data Dupont et al. (2022) to classifying and editing implicit neural representations De Luigi et al. (2023). With the proliferation of model repositories such as HuggingFace and CivitAI, developing methods that effectively learn directly from weights has become increasingly vital Kahana et al. (2024). To date, the literature has predominantly focused on utilizing these weights as inputs and outputs to secondary models, leaving their potential as intermediate representations (e.g. latent vectors, hidden states) in end-to-end training systems unexplored.
@@
Concurrently, linear Recurrent Neural Networks (RNNs) have experienced a significant resurgence, driven by their hardware-friendly nature and the resulting ease of training Dao and Gu (2024). Linearity facilitates advanced parallelization capabilities Movahedi et al. (2025), and has enabled exceptional performance on long-sequence tasks Gu et al. (2021); Orvieto et al. (2023). That said, recent findings question the information content of their compressed state representations Merrill et al. (2024). Additionally, an important body of work establishes how linear Transformers Vaswani et al. (2017) and a particular instantiation of linear RNNs — State-Space Models (SSMs) — are fundamentally less powerful than standard RNNs Beck et al. (2024); Deletang et al. (2023); Merrill and Sabharwal (2023). These strongly suggest that non-linearities are essential for the expressivity of deep sequence models.
@@
The preceding analyses motivate our examination of gradient-free adaptation through the lens of weight-space linear RNNs. To harness the strengths of the aforementioned paradigms, we formulate several research questions: $\bullet$ Can the weights of an auxiliary (root) neural network serve as high-resolution hidden states for linear RNNs? $\bullet$ Can the resulting architecture adapt its dynamics during inference without requiring gradient computation? $\bullet$ Can the activation functions in the root network provide sufficient non-linearity to significantly enhance the expressive capacity of such models?
We answer all these questions in the affirmative. Specifically, our technical contributions are summarized as follows:
@@
    We propose a novel framework (WARP) for learning on sequential data, integrating weight-space features, sequence mixing via linear recurrence, and nonlinear self-decoding which enables the embedding of domain-specific inductive biases.
@@
    Following the SSM paradigm, we introduce multiple training algorithms, including an auto-regressive mode well-suited for noisy sequence. WARP demonstrates competitive or state-of-the-art performance across diverse forecasting and classification benchmarks.
@@
    We establish a comprehensive benchmark suite to evaluate various capabilities of RNNs regarding reconstruction, adaptation and memory-retention in physical systems. WARP consistently outperforms traditional RNNs, SSMs, and Transformer architectures.
@@ 2 Weight-space Adaptive Recurrent Prediction (WARP)
State Transition & Decoding in Recurrent Neural Networks Standard RNNs $\mathbf{h}_{t}=f_{\Phi}(\mathbf{h}_{t-1},\mathbf{x}_{t})$ $\mathbf{y}_{t}=g_{\Psi}(\mathbf{h}_{t})$ Linear RNNs $\mathbf{h}_{t}=A\mathbf{h}_{t-1}+B\mathbf{x}_{t}$ $\mathbf{y}_{t}=C\mathbf{h}_{t-1}$ Weight-Space Linear RNNs $\theta_{t}=A\theta_{t-1}+B(\mathbf{x}_{t}-\mathbf{x}_{t-1})$ $\mathbf{y}_{t}=\theta_{t}(\tau)$
Figure 1: Conceptual comparison between various RNN architectures. Standard RNNs (e.g. Hochreiter and Schmidhuber (1997)) feature a non-linear transition function $f_{\Phi}$ unlike their linear counterparts (e.g. Orvieto et al. (2023)). Our proposed WARP is a Weight-Space Linear RNN, with hidden states represented as $\theta_{t}$ to emphasize their view as weights of a self-decoding function (in addition to their view as tensors). In the decoding phase, $\tau=t/(T-1)$ with $T$ the training sequence length.
This section introduces WARP, a novel framework for deep sequence models that operate by directly modulating the weights of a function approximator in response to sequential inputs. We begin by establishing the problem setting, followed by WARP’s architectural and training details.
@@ 2.1 Problem Setting
Our framework addresses the canonical sequence modeling problem, wherein a computational model must establish a mapping from an input $\mathbf{x}_{t}\in\mathbb{R}^{D_{x}}$ to a corresponding output $\mathbf{y}_{t}\in\mathbb{R}^{D_{y}}$, with $t\in\{0,\ldots,T-1\}$ denoting the time step index. The integer $T>0$ represents the training sequence length, which remains invariant across all sequences within a training batch [^2] We note that $T$ may be different for testing sequences. $D_{\bullet}$ is the dimensionality of the sub-scripted quantity.. We assume that all input sequences are sampled at the same uniform intervals. With $B$ representing the batch dimension, our models establish a mapping from $\mathbb{R}^{B\times T\times D_{x}}$ to $\mathbb{R}^{B\times T\times D_{y}}$.
In the regression setting of time series forecasting, we have $\mathbf{y}_{t}\triangleq\mathbf{x}_{t+1}$, as our objective is to predict future tokens conditioned on a preceding sequence of tokens, designated as the “context” $\mathbf{x}_{<L}\triangleq\{\mathbf{x}_{t}\}_{t\in\{0,\ldots,L-1\}}$, where $L$ denotes the context length. Consistent with the sequence modeling literature, we desire the ability to perform auto-regressive rollouts during inference. For classification tasks, only the final token $\mathbf{y}_{T-1}$ is treated as a softmax-activated logit to assign a label to the sequence.
@@ 2.2 Architecture
Unlike traditional recurrent networks that update hidden states $\mathbf{h}_{t},\forall\,t\in\{0,\ldots,T-1\}$, weight-space linear RNNs such as WARP update the flattened parameters of a root neural network $\theta_{t}$, effectively learning a dynamics model in weight-space (see Figure 1). Specifically, we define the recurrence relation and the subsequent decoding:
$\theta_{t}=A\theta_{t-1}+B\Delta\mathbf{x}_{t},\qquad\text{and}\qquad\mathbf{y}_{t}=\theta_{t}(\tau),$ (1)
where $\theta_{t}\in\mathbb{R}^{D_{\theta}}$ represents the weights of the root neural network at time step $t$, and $\Delta\mathbf{x}_{t}=\mathbf{x}_{t}-\mathbf{x}_{t-1}$ is the input difference. $A\in\mathbb{R}^{D_{\theta}\times D_{\theta}}$ is the state transition “weights-to-weights” matrix, and $B\in\mathbb{R}^{D_{\theta}\times D_{x}}$ the input transition “data-to-weights” matrix. Both $A$ and $B$ are learned during the training process defined below. To compute the output $\mathbf{y}_{t}$, the (unflattened) hidden state $\theta_{t}(\cdot)$ is viewed as a non-linear decoding function applied to the normalized training time $\tau=t/(T-1)$. Compared to other RNNs, $\theta_{t}$ plays both the roles of the hidden state and the decoder, effectively decoding itself. This self-decoding significantly saves on learnable parameter count.
@@
Importantly, all hidden states can be computed efficiently thanks to the linear recurrence in Equation 1. Once materialized, all $\theta_{t}$s can be decoded in parallel. This allows our model to combine the efficiency of linear recurrence with the expressivity enabled by incorporating non-linearities.
@@
Another key aspect of our formulation is the use of input differences $\Delta\mathbf{x}_{t}$ rather than direct inputs $\mathbf{x}_{t}$, which is a choice Kidger et al. (2020) theoretically motivate for continuous-time RNNs. When inputs change slowly or remain constant, the weight updates become proportionally smaller, and vice-versa. WARP essentially learns to convert input inferences into neural network updates, a critical self-supervision ability for continual learning and test-time adaptation Behrouz et al. (2024).
Root network.
@@
The root network $\theta_{t}$ is implemented as a multilayer perceptron (MLP) McCulloch and Pitts (1943) with a 1-dimensional input $\tau\in\mathbb{R}^{1}$, and output dimension either $D_{y}$ or $2\times D_{y}$ depending on whether uncertainty estimation is performed. In such settings, the network predicts in addition to a mean $\hat{\boldsymbol{\mu}}_{t}\in\mathbb{R}^{D_{y}}$, a quantity $\tilde{\boldsymbol{\sigma}}_{t}\in\mathbb{R}^{D_{y}}$ on which a positivity-enforcing function is applied to obtain an element-wise standard deviation:
$\hat{\boldsymbol{\sigma}}_{t}=\max(\text{softplus}(\tilde{\boldsymbol{\sigma}}_{t}),\sigma_{\min}),$ (2)
where $\sigma_{\min}$ is a fixed positive problem-dependent lower bound for numerical stability.
Initialization.
@@
Similar to prior work Le et al. (2015), the state transition matrix $A$ is initialized as the identity operator $I_{D_{\theta}\times D_{\theta}}$. This emulates gradient descent and residual connections in ResNets He et al. (2016), thereby facilitating gradient flow during backpropagation through time Werbos (1990). We find that initializing the input transition matrix $B$ as the zero matrix $\mathbf{0}_{D_{\theta}\times D_{x}}$ is useful to ensure that the sequence of weights $\theta_{t}$ does not diverge early on in the training. This strategic initialization also imposes a critical constraints wherein the initial hidden state $\theta_{0}$ must encode semantically rich information applicable to the entire sequence.
@@
The initial weights $\theta_{0}$ are determined by processing the first observation: $\theta_{0}=\phi(\mathbf{x}_{0})$, where the “initial network” $\phi$ is a hypernetwork Ha et al. (2016) defined as a learnable MLP with gradually increasing width. On sequence modeling problems with fixed or mildly-varying initial conditions, we sidestep $\phi$ and directly learn $\theta_{0}$, which is initialized with classical techniques Glorot and Bengio (2010); He et al. (2015) (and subsequently flattened into a 1D vector).

### v2

@@ Abstract
We introduce WARP (Weight-space Adaptive Recurrent Prediction), a simple yet powerful model that unifies weight-space learning with linear recurrence to redefine sequence modeling. Unlike conventional recurrent neural networks (RNNs) which collapse temporal dynamics into fixed-dimensional hidden states, WARP explicitly parametrizes its hidden state as the weights and biases of a distinct auxiliary neural network, and uses input differences to drive its recurrence. This brain-inspired formulation enables efficient gradient-free adaptation of the auxiliary network at test-time, in-context learning abilities, and seamless integration of domain-specific physical priors. Empirical validation shows that WARP matches or surpasses state-of-the-art baselines on diverse classification tasks, featuring in the top three in 5 out of 6 real-world challenging datasets. Furthermore, extensive experiments across sequential image completion, multivariate time series forecasting, and dynamical system reconstruction demonstrate its expressiveness and generalisation capabilities. Remarkably, a physics-informed variant of our model outperforms the next best model by more than 10x. Ablation studies confirm the architectural necessity of key components, solidifying weight-space linear RNNs as a transformative paradigm for adaptive machine intelligence.
@@ 1 Introduction
Deep sequence models, which continuously drive progress in machine learning, are limited in their ability to operate outside their training distribution [2; 39; 32]. For instance, subsets of Neural ODE parameters [17] necessitate adaptation via gradient descent to maintain performance on out-of-distribution (OoD) sequences [48; 73]. While effective, their explicit gradient calculation cost has recently catalysed research into gradient-free test-time adaptation methods [83; 67; 42]. This surge of interest is embodied by in-context learning [59; 83], which has recently been shown to perform test-time adaptation since during inference, it implicitly minimises a loss objective using gradient information [91; 99]. Another reason for the poor generalisation of discrete deep sequence models is the inability to inject domain-specific priors in their forward pass. In an effort to preserve all desirable traits while unleashing a breadth of possibilities, we combine two of the most powerful emerging deep learning paradigms: weight-space learning and linear recurrence.
@@
Weight-space learning — the paradigm that treats the weights and biases of a function approximator as data points for another learning system [82] — offers unprecedented potential for extracting properties of a trained model solely from its “weights” [^1] Following the convention from [103], we refer to the learnable parameters of the processed function approximator as ‘weights’ (or ‘weight space’ to indicate the space they belong to) and those of the higher-level learning system (e.g., the neural functional) as simply ‘parameters’.. Applications span from predicting generalisation error [88] and recovering training data [27] to classifying and editing implicit neural representations [23]. With the proliferation of model repositories such as HuggingFace and CivitAI, developing methods that effectively learn directly from weights has become increasingly vital [47]. To date, the literature has predominantly focused on utilizing these weights as inputs and outputs to higher-level models, leaving their potential as intermediate representations (e.g., latent vectors, hidden states) in end-to-end training systems unexplored.
State Transition & Decoding in Recurrent Neural Networks Standard RNNs $\mathbf{h}_{t}=f_{\Phi}(\mathbf{h}_{t-1},\mathbf{x}_{t})$ $\mathbf{y}_{t}=g_{\Psi}(\mathbf{h}_{t})$ Linear RNNs $\mathbf{h}_{t}=A\mathbf{h}_{t-1}+B\mathbf{x}_{t}$ $\mathbf{y}_{t}=C\mathbf{h}_{t-1}$ Weight-Space Linear RNNs $\theta_{t}=A\theta_{t-1}+B(\mathbf{x}_{t}-\mathbf{x}_{t-1})$ $\mathbf{y}_{t}=\text{MLP}_{\theta_{t}}(\tau)$
Figure 1: Background and conceptual comparison between RNN architectures. Standard RNNs (e.g. [43; 18]) feature a non-linear transition function $f_{\Phi}$ unlike their linear counterparts (e.g. [36; 75]). Our proposed weight-space linear RNNs view their hidden state — denoted as $\theta_{t}$ — as the parameters of a family of functions. As observed in the bottom-right corner, $\theta_{t}$ represents, in the general case, the flattened weights of an MLP at time step $t$. Its input $\tau$ is a (concatenation of) coordinate system(s) to maximally make use of the canonical ordering of the sequence.
@@
Concurrently, linear Recurrent Neural Networks (RNNs) have seen a notable resurgence, largely due to their hardware efficiency and the resulting ease of training [21]. Linearity enables advanced sequence parallelisation techniques [84; 69; 97] and has delivered exceptional performance on long-sequence tasks [36; 75]. However, recent findings raise concerns about the information capacity of their compressed state representations [65]. Moreover, a substantial body of work has shown that linear Transformers [49] and State-Space Models (SSMs) [36] — a particular instantiation of linear RNNs — are fundamentally less expressive than the standard non-linear RNNs depicted in Fig. 1 [9; 25; 64]. Taken together, these results strongly suggest that non-linearities are crucial for the expressivity of deep sequence models. They invite the reintroduction of non-linearities into sequence models, while maintaining the hardware-friendly nature of linear RNNs.
@@
The preceding analyses motivate our examination of weight-space linear RNNs. To harness the strengths of its constituting paradigms, we formulate several research questions: $\bullet$ Can the weights of an auxiliary function approximator serve as high-resolution hidden states for linear RNNs? $\bullet$ Can that auxiliary function be effectively adapted during inference without requiring gradient computation? $\bullet$ Are the non-linearities in the auxiliary function approximator sufficient to significantly enhance the expressive power of such models?
We answer these questions in the affirmative by proposing Weight-space Adaptive Recurrent Prediction (WARP) models as powerful expressions of weight-space linear RNNs, which we illustrate in Fig. 1. Specifically, our original contributions can be summarised as follows:
@@
    We formulate a general framework for sequence modelling in weight-space, blending linear recurrence with non-linear decoding. Rather than relying on direct inputs, we draw inspiration from the human brain and compute signal differences to drive such recurrences. To the best of our knowledge, our framework is the first of its kind to treat weight-space features as intermediate hidden state representations in a recurrence.
@@
    To train weight-space linear RNNs, we introduce two parallelisable algorithms: a convolutional mode and an efficient recurrent mode (with and without support for auto-regression) well-suited for noisy sequences. These algorithms unlock three practical use cases: ($i$) gradient-free adaptation, i.e., the ability to update model behaviour while scanning a sequence without ever explicitly computing gradients; ($ii$) in-context learning, i.e., the ability to perform predictions solely based on information present in the sequence’s context; and (iiiiii) physics-informed modelling, i.e., the ability to incorporate domain-specific continuous physical priors in the discrete linear recurrence. This final core application is evidenced in our WARP-Phys model, which achieves an order of magnitude lower error over WARP on a wide set of synthetic dynamical system reconstruction datasets.
@@
    We identify an extensive suite of real-world benchmarks to evaluate various capabilities of RNNs regarding classification, reconstruction, adaptation, and memory retention. Empirical results demonstrate how WARP consistently matches or outperforms traditional RNNs, SSMs, and Transformer architectures. Remarkably, we push the state-of-the-art by featuring in the top three in 5 out of 6 multivariate time series classification datasets necessitating the understanding of both short- and extremely long-range dependencies.
@@ 2 Weight-space Adaptive Recurrent Prediction (WARP)
This section presents the core ideas underpinning weight-space linear recurrence, our novel framework for deep sequence modelling that operates by directly modulating, in response to sequential input differences, the weights of a function approximator [6]. Out of simplicity and consistency with the related literature in Appendix A, we focus in the remainder of this paper on the WARP model, which modulates a feed-forward neural network [63]. We begin by establishing the problem setting, followed by WARP’s architectural and training details.
@@ 2.1 Problem Setting
Our framework addresses the general sequence modelling problem, wherein a computational model must establish a mapping from an input $\mathbf{x}_{t}\in\mathbb{R}^{D_{x}}$ to a corresponding output $\mathbf{y}_{t}\in\mathbb{R}^{D_{y}}$, with $t\in\{0,\ldots,T-1\}$ denoting the time step index. The integer $T>0$ represents the training sequence length, which remains invariant across all sequences within a training batch [^2] We note that $T$ may be different for testing sequences. $D_{\bullet}$ is the dimensionality of the subscripted quantity.. We assume that all input sequences are sampled at the same uniform intervals. Ignoring the batch dimension for simplicity, our models establish a mapping from $\mathbb{R}^{T\times D_{x}}$ to $\mathbb{R}^{T\times D_{y}}$ (see Fig. 2).
Figure 2: (Left) General sequence modelling setting. In the forecasting scenario, for instance, a context of length $L$ informs the prediction of future states. (Right) WARP’s unfolded recurrence. The initial hypernetwork $\phi$ and transition matrices $(A,B)$ — highlighted in orange — are learnable parameters, fitted via conventional gradient descent.
In the regression setting of time series forecasting, we have $\mathbf{y}_{t}\triangleq\mathbf{x}_{t+1}$, as our objective is to predict future tokens conditioned on a preceding sequence of tokens, designated as the “context” $\mathbf{x}_{<L}\triangleq\{\mathbf{x}_{t}\}_{t\in\{0,\ldots,L-1\}}$, where $L$ denotes the context length. Critically, we desire the ability to perform auto-regressive rollouts during inference. For classification tasks, only the final token $\mathbf{y}_{T-1}$ is treated as a softmax-activated logit to assign a label to the sequence.
@@ 2.2 Architecture
While traditional recurrent networks update obscure hidden states $\mathbf{h}_{t},\forall\,t\in\{1,\ldots,T-1\}$, weight-space linear RNNs such as WARP update the weights and biases of an auxiliary “root” neural network $\theta_{t}$, effectively learning a dynamics model in weight-space (see Figs. 1 and 2). Specifically, we define the recurrence relation and the subsequent decoding:
$\theta_{t}=A\theta_{t-1}+B\Delta\mathbf{x}_{t},\qquad\text{and}\qquad\mathbf{y}_{t}=\text{MLP}_{\theta_{t}}(\tau),$ (1)
where the hidden state $\theta_{t}\in\mathbb{R}^{D_{\theta}}$ represents the flattened weights of the root neural network at time step $t$, and $\Delta\mathbf{x}_{t}=\mathbf{x}_{t}-\mathbf{x}_{t-1}$ is the input difference. $A\in\mathbb{R}^{D_{\theta}\times D_{\theta}}$ is the state transition “weights-to-weights” matrix, and $B\in\mathbb{R}^{D_{\theta}\times D_{x}}$ the input transition “data-to-weights” matrix. To compute the output $\mathbf{y}_{t}$, the vector $\theta_{t}$ is unflattened and combined with non-linear static activation functions to reconstitute the MLP root network. This decoding function is applied to $\tau$, a coordinate system (or a concatenation thereof) that suitably informs the model of the canonical ordering of the sequences at hand. Powerful examples of coordinate systems (see Section B.2.1) include normalised pixel locations (for images viewed as sequences), normalised training time $\tau=t/(T-1)$, or the general positional encoding to facilitate generalisation beyond $T$ [89].
Compared to other RNNs, $\theta_{t}$ plays both the roles of the hidden state and the parameters of the decoder, effectively decoding itself. Such self-decoding significantly saves on learnable parameter count.
@@
Importantly, all hidden states can be precomputed efficiently thanks to the linear recurrence in Eq. 1, using for instance, the parallel “scan” operator [84]. Once materialised, the $\theta_{t}$ can be reconstituted and self-decoded independently. This allows our model to combine the efficiency of linear recurrence with the expressivity enabled by incorporating non-linearities.
@@
Another key aspect of our formulation is the use of input differences $\Delta\mathbf{x}_{t}$ rather than direct inputs $\mathbf{x}_{t}$, which is a choice Kidger et al. [53] theoretically motivated for continuous-time RNNs. When inputs change slowly or remain constant, the weight updates become proportionally smaller, and vice-versa. WARP essentially learns to convert input differences into neural network updates, a critical self-supervision ability for continual learning and test-time adaptation [10].
Architecture of the root network.
@@
The root network $\theta_{t}$ is implemented as a fixed-width multilayer perceptron (MLP) [63] with a $D_{\tau}$-dimensional input, and output dimension either $D_{y}$ or $2\times D_{y}$ depending on whether uncertainty measures are required in the pipeline. When modelling uncertainty, the network predicts in addition to a mean $\hat{\bm{\mu}}_{t}\in\mathbb{R}^{D_{y}}$, a quantity $\tilde{\bm{\sigma}}_{t}\in\mathbb{R}^{D_{y}}$ on which a positivity-enforcing function is applied to obtain an element-wise standard deviation $\hat{\bm{\sigma}}_{t}=\max(\text{softplus}(\tilde{\bm{\sigma}}_{t}),\sigma_{\min}),$ where $\sigma_{\min}$ is a fixed positive problem-dependent lower bound for numerical stability.
Initialisation of learnable parameters.
@@
Similar to prior work [57], the state transition matrix $A$ is initialised as the identity operator $I_{D_{\theta}\times D_{\theta}}$. This emulates gradient descent and residual connections in ResNets [41], thereby facilitating gradient flow during backpropagation through time [95]. We find that initializing the input transition matrix $B$ as the zero matrix $\mathbf{0}_{D_{\theta}\times D_{x}}$ is useful to ensure that the sequence of weights $\theta_{t}$ does not diverge early on in the training. This strategic initialisation also imposes a critical constraint wherein the initial hidden state $\theta_{0}$ must encode semantically rich information applicable to the entire sequence.
@@
The initial weights $\theta_{0}$ are determined by processing the first observation: $\theta_{0}=\phi(\mathbf{x}_{0})$, where the “initial network” $\phi$ is a hypernetwork [38] defined as a learnable MLP with gradually increasing width (see Fig. 2). On sequence modelling problems with fixed or mildly-varying initial conditions, we sidestep $\phi$ and directly learn $\theta_{0}$, which is initialised with classical techniques such as Glorot [29] or He [40] (and subsequently flattened into a 1D vector).